OpenStudy (anonymous):

What is the solution to the equation? 4(x – 3) = 7x + 24 A. x = –36 B. x = 12 C. x = 4 D. x = –12 is it D?

OpenStudy (anonymous):

@Sophie_dofi

OpenStudy (anonymous):

Yes, it's -12, is there a reason why you're unsure about it?

OpenStudy (anonymous):

i dont know how to solve it

OpenStudy (anonymous):

here lets do it step by step ok?

OpenStudy (anonymous):

we want to get x moved to one side of the equation

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

4(x-3) = 7x + 24 4x - 12 = 7x +24 4x - 7x = 24 + 12 -3x = 36 3x = -36 x = -12

OpenStudy (anonymous):

check: 4[(-12)-3)] = 7(-12) +24 4[(-15)]= (-84) +(24) 4(-15) = -60 and (-84)+ (24) = -60

OpenStudy (anonymous):

This is solved: x = -12